# Forest-Chat: Adapting Vision-Language Agents for Interactive Forest Change Analysis

### v1

@@ 5.6 Forest-Chat as a Tool for Forest Change Analysis
To qualitatively assess benchmark model performance on the Forest-Change dataset, Figure 6 presents representative examples of change mask and change caption predictions from their respective task capable models. For change mask generation, all change detection capable methods are capable of capturing larger and medium change areas, with the FC-Supervised the most accurate for smaller change regions. All models struggle with accurately predicting numerous small, scattered change regions. When used in a zero-shot capacity, Forest-Chat is notably more susceptible to producing false positives (red), with extreme sensitivity to atmospheric noise. Like the supervised methods, it is capable of detecting the main regions and patterns of change, but often has little overlap when encountering small scattered loss patterns. Smoother change boundaries and spatial coherence are observed for both supervised and zero-shot settings in Forest-Chat compared to the other benchmarked methods, likely due to attention-based feature aggregation and region-level decoding.
For generating change captions describing forest change, only FC-Supervised is able to construct sentences that reliably include semantic information denoting change severity, location, and patch characteristics. Of these descriptors, it is most consistent with severity. When describing the locality and patchiness, the caption quality can be seen to occasionally degrade, such as with examples 1, and 4. Both Change3D and Chg2Cap routinely exhibit issues with providing sufficient semantic detail and well-formed sentence compositions. Due to training performances optimising against BLEU-4, all models have learnt to replicate the rule-based generation inherent to the majority of the captions within the training dataset, almost entirely ignoring the semantic language provided by the human captions. Generally, the caption quality is encouraging, but all models will require stronger training data and domain-aware learning to improve lexical richness and geographic domain grounding.
@@
Figure 7 presents two representative conversations with the Forest-Chat agent, which incorporates the supervised change interpretation, zero-shot change detection, additional tooling, and a ChatGPT-4o-mini instance for task orchestration and interaction. The conversations highlight the agent’s ability to respond to user queries through natural language and provide appropriate information according to the request. Demonstrated within the conversation is the ability to produce change masks via supervised and zero-shot methods, provide descriptive change captions, estimate the percentage of area affected by deforestation, count the number of cleared patches, and reason about future change. The agent provides this rich degree of interaction and analysis through a combination of its innate knowledge and the toolset provided to it.
In line with claims made by Xu et al. (2024), the overall toolset available to the agent is small, and new functionality needs to be manually incorporated into the agent, with few-shot examples helping to guide its awareness of how to utilise the tool. Additionally, the agent would periodically fail to construct valid executable Python code, or get confused about how to respond to the user’s request. This can occur due to the limited support that few-shot examples provide, since they may not adequately cover an expanding range of supported tasks. Providing a task-aware retrieval strategy that can access predefined task solutions, can significantly improve the utilisation of toolsets (Xu et al., 2024).
Figure 7: Conversation examples between users and Forest-Chat. Forest-Chat is capable of understanding users and orchestrating the relevant tools to answer questions relating to forest change detection, building upon the capabilities of Change-Agent.
@@ 6 Discussion and Future Work
The development of Forest-Chat demonstrates the feasibility of adapting VLMs with conversational interfaces for forest change analysis. The system can generate both pixel-level change masks and semantic captions that allow users to explore temporal dynamics interactively. The system effectively identifies dominant and visually salient change patterns, demonstrating the value of aligning visual interpretation with natural language interaction. At the same time, the results highlight two important remaining challenges: detecting small, spatially fragmented forest change patches, and identifying visually subtle but ecologically significant changes.
@@
Compared with other datasets for RSCDC, Forest-Change exhibits notable limitations in terms of scale and diversity. The RS imagery and change masks are derived from a limited number of unique sites (Hewarathna et al., 2024). Additionally, the generation of the majority of captions via a programmatic approach compounded by the limited mask diversity resulted in suboptimal caption quality and an artificial limitation in linguistic complexity that largely ignores each scene’s geographical context. Developing a large, diverse forest change dataset with high-quality annotations is therefore crucial, particularly to capture subtle, localised, and ecologically meanginful change processes. Integrating multiple modalities and various temporal sequence lengths would increase the variety of downstream tasks models could learn from the data.
@@
Designing agentic VLM frameworks capable of providing further support for forest change analysis tasks is another important avenue. The current framework is limited to tasks auxiliary to change detection, captioning, and general reasoning. Enhanced visualisation tools and regression outputs can be applied for advanced ecological health assessments and biodiversity monitoring. If pursued through a single multi-task architecture, appropriate model structures and training strategies will need to be considered, with importance given to task balance and inter-task relationships. Although FC-Supervised contains architectural components that improve the ability to detect small and fine-grained change objects, performance on these subtle patterns remains limited. Future work should aim to develop architectures that can reliably capture small-scale forest changes informed by ecological dynamics and domain-specific knowledge.
The reliance on few-shot examples within the system prompt also results in periodic task execution failures and is inefficient for scaling to more complex scenarios. In such cases, the LLM may misinterpret user instructions or fail to produce useful solutions. Expanding the agent’s toolset, integrating task-aware retrieval, and supporting a multi-agent framework, or Mixture-of-Experts system would markedly enhance robustness for a wide array of applications. Storing previous conversations in a knowledge base would further help to ground future LLM responses. Enabling zero-shot change captioning and embedding geographic knowledge could further ground reasoning and reduce the reliance on large labelled datasets.
@@ 7 Conclusion
Forest-Chat addresses the growing need for interpretable and accessible forest change analysis by integrating VLMs with a conversational interface. The system leverages the MCI model in combination with AnyChange to provide pixel-level change masks and semantic captions, enabling comprehensive understanding of bi-temporal forest imagery and forest loss events. The introduction of the Forest-Change dataset, which pairs change masks with descriptive captions, supports joint training for detection and captioning tasks, while multi-task learning balancing ensures effective optimisation across objectives. Supplementary tools, including a bi-temporal image captioning interface and a zero-shot point query interface, further facilitate interactive exploration and annotation, enhancing usability for researchers and practitioners.
Extensive experiments demonstrate that Forest-Chat can reliably detect and describe forest changes, with the underlying MCI model generalising well from an urban change context into a forest one, showing superior or competitive performance to SOTA models for both tasks. Incorporating an LLM with VLM capabilities lowers the barriers for engaging with RS data, highlighting the broader potential of interactive AI systems to improve interpretability, accessibility, and efficiency in environmental monitoring. This framework provides a foundation for future expansions, including scaling to larger and more diverse datasets, integrating domain-knowledge and additional modalities, and supporting richer natural language queries and tasks, advancing the development of intelligent tools for ecological research and forest monitoring.
@@ Acknowledgements
All data and code used in this work are available in the following GitHub repository: https://github.com/JamesBrockUoB/ForestChat.
@@
Appendix A Supplementary Results
Table 1: Full MTL strategy change detection and change captioning ablation results. Mean $\pm$ standard deviation across configurations, with three runs each. Bold indicates the best mean value per metric within each group when the improvement is clearly distinguishable relative to standard deviation.
@@ Appendix A Supplementary Results
Table 2: Full MTL change detection and captioning ablation results shown as percentage changes relative to baseline. Only $\Delta\%$ values are reported; positive/negative values indicate improvement or decline compared to the baseline configuration respectively. Bold indicates the largest improvement among non-baseline configurations per metric when the improvement is clearly distinguishable.
@@
Figure 1: Training Time Comparison by Backbone Size for Each Dataset. The lines show the training time for one epoch (mean over three runs) per backbone size for the LEVIR-MCI-Trees and Forest-Change datasets.

### v2

@@ 5.6.3 Forest-Chat as a Tool for Forest Change Analysis
To qualitatively assess benchmark model performance on the Forest-Change dataset, Figure 7 presents representative examples of change mask and change caption predictions from each model capable of the respective task. For change mask generation, all change detection capable methods are capable of capturing larger and medium change areas, with the FC-Supervised the most accurate for smaller change regions. All models struggle with accurately predicting numerous small, scattered change regions. This is likely attributable to the underrepresentation of small-patch examples during training, where the loss signal is dominated by the larger no-change class and small changed objects contribute disproportionately few pixels to overall mIoU. Incorporating auxiliary data such as canopy height maps could provide additional feature channels to help distinguish small change regions from surrounding vegetation, while architectural components designed for fine-grained localisation and boundary refinement represent a further avenue for mitigating these failure modes. When used in a zero-shot capacity, Forest-Chat is notably more susceptible to producing false positives, with extreme sensitivity to atmospheric noise, and often has little overlap when encountering small scattered loss patterns. Smoother change boundaries and spatial coherence are nonetheless observed for both supervised and zero-shot settings compared to the benchmarked methods, likely due to attention-based feature aggregation and region-level decoding.
For change captioning, both FC-Supervised and FC-Zero-shot reliably construct sentences capturing change severity, location, and patch characteristics. FC-Supervised is most consistent in capturing severity, though quality can occasionally degrade when describing locality and patchiness, as seen in examples 1 and 4. FC-Zero-shot constructs semantically valuable descriptions for both prompt types, but is considerably more aligned with target captions under style-guided prompting. Both Change3D and Chg2Cap routinely exhibit issues with semantic detail and sentence composition. Due to training objectives optimising against BLEU-4, all supervised models tend to replicate the rule-based phrasing present in the majority of training captions, largely ignoring the richer semantic language of human-written examples. All models are constrained by the limited linguistic diversity of training captions; richer, geographically grounded annotations would directly improve both lexical variety and spatial specificity. Common failure cases can be partially resolved through FC-Zero-shot refinement, which enriches captions with domain-aware semantics, although supervised-level performance is not reached even with style-guided prompting.
@@
Figure 8 presents two representative conversations with the Forest-Chat agent, demonstrating its ability to respond to user queries through natural language across a range of tasks including supervised and zero-shot change detection, change captioning, deforestation area estimation, patch counting, and future change reasoning. The agent delivers this range of capabilities through a combination of its innate knowledge and the provided toolset.
The overall toolset available to the agent is intentionally compact, with new functionality requiring manual incorporation alongside few-shot examples to guide tool usage awareness. Although the agent periodically fails to construct valid executable code or misunderstands aspects of a user’s request, internal validation and retry mechanisms prevent erroneous responses from reaching the user. Logic failures can occur when few-shot examples do not adequately document the range of supported tasks. Exploring prompt engineering optimisation or task-aware retrieval strategies that can access pre-defined task solutions and forest ecology knowledge could significantly improve toolset utilisation and conversational richness (Xu et al., 2024).
Figure 8: Conversation examples between users and Forest-Chat. Forest-Chat is capable of understanding users and orchestrating relevant tools to answer questions relating to forest change interpretation.
@@ 6 Discussion and Future Work
The development of Forest-Chat demonstrates the feasibility of adapting VLMs with conversational interfaces for forest change analysis. The system can generate both pixel-level change masks and semantic captions that allow users to explore temporal dynamics interactively, effectively identifying dominant and visually salient change patterns. At the same time, the results highlight two important remaining challenges: detecting small, spatially fragmented forest change patches, and identifying visually subtle but ecologically significant changes. From a practical deployment perspective, FC-Supervised’s inference times as reported in Table 2 are not prohibitive, with training feasible on consumer-grade CPU hardware for smaller dataset sizes. The system’s end-to-end query response times are well-suited to an analytical workflow where interpretability takes precedence over real-time interaction. Where inference efficiency is a priority, training FC-Supervised or fine-tuning from the provided pretrained weights offers a more computationally efficient and higher-performing alternative to FC-Zero-shot change detection, as cross-domain transfer experiments in Section 5.5 demonstrate that only limited domain samples are required. Taken together, these findings underscore the broader challenge of deploying RSICI systems beyond urban and general-purpose domains, with Forest-Chat serving as an empirical study of the considerations involved in such adaptation.
@@
Compared with other datasets for RSCDC, Forest-Change exhibits notable limitations in terms of scale and diversity. The RS imagery and change masks are derived from a limited number of unique sites (Hewarathna et al., 2024), and domain transfer experiments on JL1-CD-Trees further highlight the challenges of generalising across datasets with substantial resolution, seasonal, and sensor differences. Additionally, the generation of the majority of captions via a programmatic approach, compounded by limited mask diversity, resulted in suboptimal caption quality. This introduced an artificial limitation in linguistic complexity, with captions largely ignoring each scene’s geographical context. We partially alleviate these issues through zero-shot caption refinement, which can enrich captions with domain-aware semantics and spatial grounding. Nonetheless, developing a large, spatially dense, and diverse forest change dataset with high-quality annotations remains crucial, particularly for capturing subtle, localised, and ecologically meaningful change processes. Integrating multiple modalities and various temporal sequence lengths would increase the variety of downstream tasks models could learn from the data.
@@
Designing agentic VLM frameworks capable of providing further support for forest change analysis tasks is another important avenue. The current framework is limited to tasks auxiliary to change detection, captioning, and general reasoning. Enhanced visualisation tools and regression outputs could be applied for advanced ecological health assessments and biodiversity monitoring. If pursued through a single multi-task architecture, appropriate model structures and training strategies would need careful consideration of task balance and inter-task relationships. Future work should aim to develop architectures that can reliably capture small-scale forest changes informed by ecological dynamics and domain-specific knowledge.
The reliance on prompt engineering with few-shot examples for agent execution logic periodically results in the LLM misinterpreting user instructions, subsequently failing to produce a useful solution, and is inefficient for scaling to a wide range of complex scenarios. Expanding the agent’s toolset, introducing forest ecology domain knowledge, integrating task-aware retrieval based on historically successful responses, and supporting a multi-agent framework or Mixture-of-Experts system would markedly enhance robustness across a broader range of operational scenarios. Systematic evaluation methodologies for LLM-driven RS agents remain an open research challenge, with recent work such as ThinkGeo (Shabbir et al., 2025) beginning to explore benchmarking frameworks for geospatial LLM system performance and capabilities.
@@ 7 Conclusion
Forest-Chat addresses the growing need for interpretable and accessible forest change analysis by integrating VLMs and MLLMs with a conversational interface. The system leverages the MCI model alongside AnyChange and GPT-4o to provide both supervised and zero-shot pixel-level change detection and semantic captioning, enabling comprehensive bi-temporal forest change analysis. The Forest-Change dataset, pairing change masks with descriptive captions, supports joint training for detection and captioning, while multi-task loss balancing ensures effective optimisation across objectives. Supplementary tools, including a bi-temporal image captioning interface and a zero-shot point query interface, further facilitate interactive exploration and annotation, enhancing usability for researchers and practitioners.
Extensive experiments demonstrate that Forest-Chat reliably detects and describes forest changes, with the MCI model generalising well from urban to forest change contexts. Cross-domain experiments on JL1-CD-Trees highlight the challenges of transferring models across datasets with substantial resolution and seasonal differences, while demonstrating that modest target-domain supervision is sufficient to achieve effective adaptation to new forest environments. Zero-shot MLLM captioning experiments show that general prompting alone is insufficient for producing domain-consistent change descriptions, although style-guided prompting improves structural consistency. Targeted refinement can correct specific captioning errors and inject missing contextual detail as a post-processing step for supervised captioning models.
Beyond system-level contributions, this work presents an empirical study of the challenges and considerations involved in adapting VLM-based agents to forest change analysis - a domain whose ecological diversity and lack of annotated data present non-trivial barriers beyond those encountered in urban RS tasks. Incorporating an LLM with vision-language capabilities lowers the barriers for engaging with RS data, highlighting the broader potential of interactive AI systems to improve interpretability, accessibility, and efficiency in environmental monitoring. This framework provides a foundation for future work, including scaling to larger and more diverse datasets, integrating domain knowledge and additional modalities, and supporting richer natural language queries and tasks. Together, these directions advance the development of intelligent, interpretable tools for ecological research and large-scale forest monitoring.
@@ Acknowledgements
All data and code used in this work are available in the following GitHub repository: ForestChat.
@@
Appendix A FC-Zero-shot Change Captioning
Prompts
All prompting modes share a common system instruction defining the model as an expert remote sensing analyst tasked with describing only genuine changes between two temporally separated aerial images.
General Prompt
The dataset-agnostic prompt instructs the model to produce a single concise sentence describing only observed changes between Image A (before) and Image B (after), or to output “the scene is the same as before” when no change is present. The prompt also enforces constraints such as avoiding references to the image identifiers and not describing unchanged content.
Style-guided Prompts
For datasets with characteristic caption distributions (Forest-Change and LEVIR-MCI-Trees), prompts additionally include few-shot examples and explicit stylistic rules to match the structure and vocabulary of the target dataset captions. These prompts specify:
  • 1.
    the allowed change vocabulary (e.g., forest loss, buildings, roads, trees),
  • 2.
    location phrasing using compound spatial descriptors (e.g., top-center, bottom-left),
  • 3.
    sentence length and grammatical style,
  • 4.
    formatting constraints specific to each dataset.
Forest-Change prompts emphasise severity descriptors and patch-based descriptions of forest loss, while LEVIR-MCI-Trees prompts prioritise object-level urban changes such as building construction or road appearance.
Refinement Prompts
Refinement operates as a two-stage pipeline. A supervised caption produced by FC-Supervised is provided to FC-Zero-shot alongside the image pair. The prompt instructs the model to preserve the original change semantics while enriching the caption with spatial grounding and additional context observable in the imagery.
Table 1: Selected FC-Zero-shot refinement examples on Forest-Change. Predicted captions are produced by FC-Supervised; refined captions are generated by FC-Zero-shot given the predicted caption and bi-temporal image pair.
Predicted Caption Refined Caption Enrichment Type
primarily occurring in the top-center and top-left forest loss is detected found scattered in numerous small patches which are highly varied in size Primarily occurring in the top-center and top-left regions, forest loss is detected found scattered in numerous small patches which are highly varied in size, with some areas showing sharp rectangular boundaries consistent with clearcutting Change process inference
mainly located across the top-center and top-left regions some modest forest loss is visible Mainly located across the top-center and top-left regions some modest forest loss is visible occurring in many small patches which are moderately varied in size with exposed understorey Forest cover context
slight forest loss is noted mainly located across the middle-right and middle-right and middle-right sections found scattered in size mainly located across the middle-right and bottom Slight forest loss is noted mainly located across the middle-right and bottom sections occurring in many small patches which are highly varied in size with exposed understorey Spatial grounding & artefact correction
primarily occurring in the middle-left and middle-left areas slight forest loss is detected occurring in many small patches which are displaying large variations in size Primarily occurring in the middle-left and center areas slight forest loss is detected occurring in many small patches which are displaying large variations in size with some fragmentation along existing clearings Geographic feature reference
Table 2: Selected FC-Zero-shot caption refinement examples on LEVIR-MCI-Trees. Predicted captions are produced by FC-Supervised; refined captions are generated by FC-Zero-shot given the predicted caption and bi-temporal image pair.
Predicted Caption Refined Caption Enrichment Type
a road and a road with villas built along the road Some villas are built along the road in the center, replacing the dense vegetation and sparse trees . Spatial localisation & land cover context
a small house appears in the bareland A small house appears near the bottom left, and a pond emerges at the top right, replacing some vegetation . Secondary change detection
a road with houses are built along the road Some houses are built around a cul-de-sac and replace the sparse vegetation alongside the road . Infrastructure reasoning
a house appears in the woods A house appears beside the existing house, and some trees are removed alongside the road . Contextual spatial grounding
Appendix B Supplementary Results
Table 1: Full MTL strategy change detection and change captioning ablation results on the LEVIR-MCI-Trees dataset. Mean $\pm$ standard deviation across configurations, with three runs each. Bold indicates the best mean value per metric within each method group. When results are tied or close, preference is given to the method with the lower standard deviation.
@@ Appendix B Supplementary Results
Table 2: Full MTL change detection and captioning ablation results on the LEVIR-MCI-Trees dataset (percentage changes relative to baseline). Only $\Delta\%$ values are reported; positive/negative values indicate improvement or decline compared to the baseline configuration respectively. Bold indicates the largest improvement among non-baseline configurations per metric when the improvement is clearly distinguishable.
Baseline (Equal) - - - - - - - - - -
+ CAGrad +0.03 +0.00 +0.08 -1.47 -2.86 -4.21 -5.23 +0.67 +1.01 +1.60
+ PCGrad +0.00 +0.01 +0.00 +0.56 -1.02 -1.86 -2.48 +3.12 +2.05 +0.32
+ GradDrop -0.26 -0.05 -0.50 -2.24 -2.83 -5.04 -7.20 -1.08 +0.28 -7.15
EDWA -0.26 -0.12 -0.41 -0.62 -1.62 -3.57 -5.82 +0.29 -0.06 -5.85
+ CAGrad -1.24 -0.40 -2.25 -6.25 -8.24 -12.84 -16.83 -6.45 -3.86 -15.81
+ PCGrad -1.38 -0.50 -2.41 -6.47 -7.84 -12.23 -14.74 -7.38 -3.61 -17.73
+ GradDrop -1.50 -0.49 -2.67 -2.75 -3.41 -5.10 -6.92 -1.57 -0.61 -5.81
Uncertainty -0.69 -0.24 -1.21 -1.18 -1.66 -1.92 -1.89 -0.15 +0.43 -1.87
+ CAGrad +0.08 +0.01 +0.19 +0.78 +0.57 +2.26 +4.86 +1.71 +1.88 +6.43
+ PCGrad -1.18 -0.43 -2.07 -2.10 -3.95 -5.48 -7.49 -0.99 +0.13 -2.34
+ GradDrop -1.03 -0.37 -1.81 -0.32 +0.00 -1.36 -2.28 +0.98 +2.00 -1.27
Table 3: Full MTL strategy change detection and change captioning ablation results on the Forest-Change dataset. Mean $\pm$ standard deviation across configurations, with three runs each. Bold indicates the best mean value per metric within each method group. When results are tied or close, preference is given to the method with the lower standard deviation.
[…]
Table 4: Full MTL change detection and captioning ablation results on the Forest-Change dataset, shown as percentage changes relative to baseline. Only $\Delta\%$ values are reported; positive/negative values indicate improvement or decline compared to the baseline configuration respectively. Bold indicates the largest improvement among non-baseline configurations per metric when the improvement is clearly distinguishable.
@@
Figure 1: Training time comparison by backbone size per dataset. The lines show the training time for one epoch (mean over three runs) per backbone size for the LEVIR-MCI-Trees and Forest-Change datasets. Segformer backbones are defined on the x-axis, whilst the training times for RegNetX-400MF are given along the y-axis.